How would a behaviourist teacher explain 1/2÷1/3?                        


1
Expert's answer
2022-05-23T18:20:20-0400

Divide the first fraction into the second fraction is the same as multiplying the first fraction by the reciprocal of the second fraction.

For example:


"\frac 1 2: \frac 1 3 = \frac 1 2 * \frac 3 1" or "1\frac 1 2"






You can also find out how many times Rectangle 2 is bigger than Rectangle 3 by counting the cells.